One leg of a right triangle is 2 more than twice the smaller leg. If the hypotenuse is 1 inch more than the larger leg, find all 3 sides.
Answer by richwmiller(17219)

b=2+2a
c=1+b
a^2+b^2=c^2
this is a classic 5,12,13 right triangle
hypotenuse c=13